# Spinning the Golden Thread: Benchmarking Long-Form Generation in Language Models

### v1

@@ Abstract
The abilities of long-context language models (LMs) are often evaluated using the ”Needle-in-a-Haystack” (NIAH) test, which comprises tasks designed to assess a model’s ability to identify specific information (”needle”) within large text sequences (”haystack”). While these benchmarks measure how well models understand long-context input sequences, they do not effectively gauge the quality of long-form text generation—a critical aspect for applications such as design proposals and creative writing. To address this gap, we have introduced a new long-form text evaluation benchmark, Spinning the Golden Thread (SGT), which tests models’ ability to identify specific events within generated long text sequences. In this benchmark, we prompt long-context LMs to create long-form text that must include particular events or constraints and evaluate their ability to incorporate these elements. We evaluated ten long-context LMs across four distinct scenarios, three types of prompt instructions, and two different generation-length settings (16K and 32K). Although these models perform well on NIAH benchmarks, none demonstrated satisfactory performance on the Spinning the Golden Thread , raising concerns about their ability to generate coherent long-form text that follows instructions. Additionally, as the length of the generated text increases, all models exhibit a significant drop in performance111We have open-sourced Spinning the Golden Thread to support the comprehensive evaluation of long-context LMs in long-form text generation. Our code and data are available at anonymousurl..
@@ 1 Introduction
This study introduces Spinning the Golden Thread (SGT), a new benchmark assessing super-long-form generation capabilities in a long-context language model. This benchmark evaluates the quality of generated content and adherence to instructions, demanding varying levels of compliance based on the specific instructions. For instance, creative writing tasks like storytelling or diary entries for fictional characters might involve detailing significant events such as family birthdays or annual vacations. Meanwhile, project planning tasks, such as urban planning, could require the inclusion of essential constraints in the document. To establish a robust benchmark, we have designed four distinct scenarios: Diary Writing, Menu Design, Skyscraper Design, and Urban Planning. Each scenario comprises three sub-scenarios with specific task instructions—single instance, range, and periodicity. These sub-scenarios capture and reflect the complexity behind long-form generation tasks, where long sequence texts must include and adhere to various criteria. These scenarios thus serve as a foundational base for evaluating long-form generation.
Our benchmark provides a fine-grained evaluation of generated text sequences by assessing for the presence of essential events or constraints specified in the prompt instructions. Unlike the NIAH (Kamradt, 2023), which tests the fundamental capability of long-context LMs to process lengthy input sequences and identify critical information, our benchmark focuses on instruction-following in super-long-form generation tasks. Specifically, it challenges models to include specific events or constraints, akin to verifying the needle in the output text sequence. This method not only captures and reflects the overall quality of the task but also minimizes the need for human evaluation or excessive reliance on LLM-based judgments. Our study findings indicate that current long-context language models are inadequate when handling SGT, even with texts shorter than 16K (see Figure 1).
@@
We introduce Spinning the Golden Thread, a dataset that includes tasks for evaluating the super-long-form generation capabilities of LLMs across multiple text token intervals (16K and 32K) and various text depth ranges.
@@ 2.2 Four Distinct Scenario Setups
In conjunction with the design of a 100-floor skyscraper, we have devised three distinct task scenarios to supplement our predefined tasks, as illustrated in Figure 2 (1). These scenarios fall into two categories: Temporal (Diary Writing, Menu Design) and Spatial (Skyscraper Design, Urban Planning). Moreover, each scenario incorporates both short and long version to assess the effectiveness of various output lengths. Each scenario is carefully engineered to evaluate the model’s capacity to produce context-sensitive and detailed lengthy texts tailored to our specific task requirements. Table 1 offers comprehensive descriptions for each scenario, with each designed around a unique template to generate 100 different task instructions.
@@ 2.3 Specific Task Instruction
For instance, in the design of a 100-floor skyscraper, the Single Instruction may specify that the 34th floor hosts an aerial gym and the 54th floor houses a law firm. The Range Instruction might designate floors 1 through 9 as a comprehensive shopping mall, whereas the Periodic Instruction could dictate that starting from the 20th floor, every 10th floor incorporates a small aerial garden. We utilize over 20 templates for each type of instruction, with the floors or locations being randomly assigned to ensure task diversity. These settings, applied via various templates, guarantee controlled coverage across all textual positions, thus facilitating a comprehensive and efficient evaluation as illustrated in Figure 2 (2). Through this approach, we generate task instructions T and simultaneously acquire the corresponding , which supports subsequent evaluations and constructs a task conducive to super-long-form generation. Subsequently, we splice the main task prompt with the specific task instructions (STI) and add the generation prompt to form the final test data.
@@
To quantitatively evaluate performance for SGT tasks, we introduce three complementary metrics:
@@
To assess the performance of specific subtasks under defined conditions, we employed the Llama3.1-8B instruction model (Dubey et al., 2024) for a few-shot evaluation. This method allowed us to verify compliance of each subtask with its corresponding Specific Task Instructions, effectively transforming the assessment into a binary classification task (Completed” and ”Uncompleted). Additionally, we manually annotated 500 data points to validate the consistency of the outcomes from the Llama3.1-8B-instruct model (Dubey et al., 2024) with human assessments. We then calculated the completion counts for the Single Instruction (SI), Range Instruction (RI), and Periodic Instruction (PI). To assess the performance of the SPI (Specific Task Instruction) process, we computed the ratio of the total number of relevant sub-scenarios included in the answers to the number of completed instructions, denoted as STIC-1
@@ 3.4 Three Specific Task Instruction
3.5 Compare with long context analysis
@@ 5 Conclusion
We present Spinning the Golden Thread SGT, a synthetic benchmark designed for evaluating the long-form generation capabilities of language models. SGT employs tasks that require following instructions over long distances, offering a comprehensive assessment of language models’ abilities. We evaluated nine advanced language models using SGT, with context sizes ranging from 32K to 128K. Although these models achieved perfect scores in the commonly used ”Ruler”, their performance declined significantly on SGT. Common failure modes observed include premature termination of the main task, incomplete responses, disregard for prior instructions as outputs lengthened, and repetitive content generation. Our findings indicate that SGT poses significant challenges for even the top-tier open-source models, highlighting considerable scope for improvement and the advantages of increasing model sizes to enhance performance in complex tasks.

### v3

# LongGenbench: Benchmarking Long-Form Generation in long-context LLMs

@@ Abstract
The abilities of long-context language models (LMs) are often evaluated using the ”Needle-in-a-Haystack” (NIAH) test, which comprises tasks designed to assess a model’s ability to identify specific information (”needle”) within large text sequences (”haystack”). While these benchmarks measure how well models understand long-context input sequences, they do not effectively gauge the quality of long-form text generation—a critical aspect for applications such as design proposals and creative writing. To address this gap, we have introduced a new long-form text evaluation benchmark, LongGenBench, which tests models’ ability to identify specific events within generated long text sequences. In this benchmark, we prompt long-context LMs to create long-form text that must include particular events or constraints and evaluate their ability to incorporate these elements. We evaluated ten long-context LMs across four distinct scenarios, three types of prompt instructions, and two different generation-length settings (16K and 32K). Although these models perform well on NIAH benchmarks, none demonstrated satisfactory performance on the LongGenBench , raising concerns about their ability to generate coherent long-form text that follows instructions. Additionally, as the length of the generated text increases, all models exhibit a significant drop in performance111We have open-sourced LongGenBench to support the comprehensive evaluation of long-context LMs in long-form text generation. Our code and data are available at https://github.com/mozhu621/LongGenBench/..
@@ 1 Introduction
This study introduces LongGenBench , a new benchmark assessing super-long-form generation capabilities in a long-context language model. This benchmark evaluates the quality of generated content and adherence to instructions, demanding varying levels of compliance based on the specific instructions. For instance, creative writing tasks like storytelling or diary entries for fictional characters might involve detailing significant events such as family birthdays or annual vacations. Meanwhile, project planning tasks, such as urban planning, could require the inclusion of essential constraints in the document. To establish a robust benchmark, we have designed four distinct scenarios: Diary Writing, Menu Design, Skyscraper Design, and Urban Planning. Each scenario comprises three sub-scenarios with specific task instructions—single instance, range, and periodicity. These sub-scenarios capture and reflect the complexity behind long-form generation tasks, where long sequence texts must include and adhere to various criteria. These scenarios thus serve as a foundational base for evaluating long-form generation.
Our benchmark provides a fine-grained evaluation of generated text sequences by assessing for the presence of essential events or constraints specified in the prompt instructions. Unlike the NIAH (Kamradt, 2023), which tests the fundamental capability of long-context LMs to process lengthy input sequences and identify critical information, our benchmark focuses on instruction-following in super-long-form generation tasks. Specifically, it challenges models to include specific events or constraints, akin to verifying the needle in the output text sequence. This method not only captures and reflects the overall quality of the task but also minimizes the need for human evaluation or excessive reliance on LLM-based judgments. Our study findings indicate that current long-context language models are inadequate when handling LongGenBench, even with texts shorter than 16K (see Figure 1).
@@
We introduce LongGenBench, a dataset that includes tasks for evaluating the super-long-form generation capabilities of LLMs across multiple text token intervals (16K and 32K) and various text depth ranges.
@@ 2.2 Four Distinct Scenario Setups
In conjunction with the design of a 100-floor skyscraper, we have devised three distinct task scenarios to supplement our predefined tasks, as illustrated in Figure 2 (1). These scenarios fall into two categories: Temporal (Diary Writing, Menu Design) and Spatial (Skyscraper Design, Urban Planning). Moreover, each scenario incorporates both short and long version to assess the effectiveness of various output lengths. Each scenario is carefully engineered to evaluate the model’s capacity to produce context-sensitive and detailed lengthy texts tailored to our specific task requirements. Table 1 offers comprehensive descriptions for each scenario, with each designed around a unique template to generate 100 different task instructions222Examples of Task Instructions for each scenario are provided in Appendix B..
@@ 2.3 Specific Task Instruction
For instance, in the design of a 100-floor skyscraper, the Single Instruction may specify that the 34th floor hosts an aerial gym and the 54th floor houses a law firm. The Range Instruction might designate floors 1 through 9 as a comprehensive shopping mall, whereas the Periodic Instruction could dictate that starting from the 20th floor, every 10th floor incorporates a small aerial garden. We utilize over 20 templates for each type of instruction, with the floors or locations being randomly assigned to ensure task diversity. These settings, applied via various templates, guarantee controlled coverage across all textual positions, thus facilitating a comprehensive and efficient evaluation as illustrated in Figure 2 (2). For instance, my wife’s birthday falls on October 8th, and our family trip to Miami is scheduled from December 24th to celebrate Christmas, among other events. Through this approach, we generate the main task instructions and simultaneously acquire the corresponding , which supports subsequent evaluations and constructs a task conducive to super-long-form generation. Subsequently, we splice the main task prompt with the specific task instructions (STI)333Each of our main task instructions splice 5 single instructions, 1 range instruction task, and 1 periodic instruction task. and add the generation prompt to form the final evalution data.
@@
To quantitatively evaluate performance for LongGenBench tasks, we introduce three complementary metrics:
@@
To assess the performance of specific subtasks under defined conditions, we employed the Llama3.1-8B instruction model (Dubey et al., 2024) for a few-shot evaluation. This method allowed us to verify compliance of each subtask with its corresponding Specific Task Instructions, effectively transforming the assessment into a binary classification task (Completed” and ”Uncompleted). Additionally, we manually annotated 500 data points to validate the consistency of the outcomes from the Llama3.1-8B-instruct model with human assessments444Detailed evaluation prompts can be found in the Appendix C.. . We then calculated the completion counts for the Single Instruction (SI), Range Instruction (RI), and Periodic Instruction (PI). To assess the performance of the SPI (Specific Task Instruction) process, we computed the ratio of the total number of relevant sub-scenarios included in the answers to the number of completed instructions, denoted as STIC-1:
@@ 3.4 Three Specific Task Instruction
3.5 Compare with long context input
@@
4 analysis and limitaion
Richness of Content
Even when sub-scenarios are specifically designed to enhance task richness, the model’s outputs tend to converge as output volume increases. This results in a homogeneity of recorded events despite differences in time and location. Such convergence not only degrades performance, but also diminishes output diversity. Attempts to mitigate this issue by adjusting the parameter during model inference have not been effective in our experiments.
Rationality of Content
Our research presently focuses only on evaluating instruction-following capabilities within tasks; a more comprehensive analysis of task rationality and integrity could be a direction for future research. For instance, if a virtual task pertains to a specific profession, all job-related activities recorded in a diary should correspond to that profession. Additionally, we noted that temperature records in diaries fail to reflect realistic temporal changes. For example, in San Francisco, one would expect temperatures to rise from cold (0-10 degrees Celsius) at the beginning of the year to warmer (20-30 degrees Celsius) by mid-year; however, our findings indicate temperatures remain in the warmer range (20-30 degrees Celsius) even into December.
Instruction Data
Significant discrepancies in the performance of most models compared to their ability to process inputs, such as Ruler(Hsieh et al., 2024), are likely due to the length distribution of instruction tuning data. These data sets are typically brief, usually under 200 tokens, and lack extended instructional content. Organizing or synthesizing instruction tuning data with longer outputs could be a potential solution for future research endeavors.
@@ 6 Conclusion
We present LongGenBench , a synthetic benchmark designed to evaluate the long-form generation capabilities of language models. LongGenBench employs tasks that require following instructions over long distances, offering a comprehensive assessment of language models’ abilities. We evaluated nine advanced language models using LongGenBench, with context sizes ranging from 32K to 128K. Although these models achieved perfect scores in the commonly used ”Ruler”, their performance decreased significantly on LongGenBench. Common failure modes observed include premature termination of the main task, incomplete responses, disregard for prior instructions as outputs lengthened, and repetitive content generation. Our findings indicate that LongGenBench poses significant challenges for even the top-tier open-source models.
@@ Appendix A Models
Appendix B FOUR DISTINCT SCENARIO SETUPS
Here are four templates for different scenarios. In Diary template, the professions and names are interchangeable options.